(i)Find the volume of the solid that is generated by revolving the region bound by the

graphs of 𝑦 = π‘₯2 and 𝑦2 = π‘₯ about the 𝑦 βˆ’ π‘Žπ‘₯𝑖𝑠.


1
Expert's answer
2022-03-24T05:22:20-0400



"V=\pi\int_0^1 ((\sqrt y)^2-(y^2)^2)dy="

"\pi\int_0^1 (y-y^4)dy=\pi(y^2/2-y^5/5)|_0^1="

"\pi(1/2-1/5)=\frac{3\pi}{10}"

Answer: "V= \frac{3\pi}{10}" .